### v1

@@ 1 Introduction
A powerful class of tactics are those which translate a goal into first order logic (FOL) or higher order logic (HOL), invoke a solver like Vampire [31], and reconstruct the proof to close the goal. These are referred to as hammers [37, 16, 28, 29, 11]. By design, these systems can only use the built-in logical connectives, cannot directly reason with dependent types, and cannot synthesize functions or objects constructively. The reconstructed proof is usually much too large to be read, and therefore is not suggested as a replacement for the tactic invocation.
In this paper, we introduce Canonical, the first solver for type inhabitation in dependent type theory. Canonical searches exhaustively for terms of any type involving dependent products ($\Pi$-types), structure types, inductive types, and let definitions. We also introduce a Lean tactic, canonical, which uses Canonical to find a proof term for the main goal. When it finds a proof term, it presents it to the user as a suggested replacement. While our implementation is not yet fully general, [^1] Specifically, we do not translate quotient types or universe levels; see Section 5. our goal is to approach soundness and completeness for the entirety of the Lean language, including definitional equality.
We demonstrate the performance of our tactic on the Natural Number Game for Lean 4 [6]. Canonical achieves an 84% on this benchmark, outperforming all existing Lean tactics. Canonical accomplishes this without any native support for natural numbers, propositional equality, logical connectives, or any other Lean definitions, and is instead provided the direct DTT representation at runtime. As such, these capabilities extend to user-defined datatypes and axiom schemas. We specifically highlight the higher-order reasoning and program synthesis capabilities of Canonical and show examples involving indexed inductive types and generalized induction.
@@ 2 Examples
Canonical is not limited to proving propositions. In this example we have an indexed inductive type Vec $\alpha$ n, representing vectors of length n with elements of type $\alpha$. Canonical is able to synthesize the append function using only the type signature, due to the enforced length constraints. The definition must be marked “noncomputable” in Lean because the code generator presently cannot compile recursors.
@@
Since Canonical searches over inhabitants of a goal type, it can also be used to find multiple inhabitants. In this example, we use Canonical to find the two endomorphisms that exist over all groups—the identity map and the constant function which returns the identity element. Using the count flag, we obtain five inhabitants with two shown above. We use a simplified group definition, Group’, to avoid reasoning about auxiliary notions like exponents.
3 Related Work
There are a number of existing theorem proving systems that are based on type inhabitation. There are sauto [15] for Rocq, Agsy [34], Mimer [48], and Auto [30] for Agda, and TWELF [47] for the Logical Framework [23]. An essential subroutine for these systems is unification, in which two terms with existentially quantified variables are made equal to each other by a substitution on those variables. An algorithm for first-order unification was described by Robinson in 1965 [45]. In the higher-order setting, unification can have multiple (or infinitely many) solutions and is undecidable [24, 21]. Nonetheless, Huet provided a procedure to enumerate higher-order unifiers in 1975 [25]. To date, no theorem provers use this algorithm on account of its inefficiency, and instead elect for decidable and efficient pattern unification [38] (or extensions thereof, e.g. [2]). In particular, the restriction to pattern unification ensures that the solutions to unification problems are unique. While this handles many common cases, this means that these systems cannot perform higher-order reasoning in general, such as the synthesis of induction predicates. Beyond this, these systems are not designed to handle the complexity of problems we are interested in solving with Canonical, such as those involving definitional equalities and requiring several nested recursors.
[…]
While these systems all have type theory support, they each sacrifice completeness or generality in some way. To our knowledge, the only account of a complete proof procedure for dependent type theory was recently given by Dowek [19]. Like Canonical, this procedure involves suspending unifications as equational constraints. However, the technique is not practical for implementation as constraints do not use the explicit substitution calculus [1, 41, 20] and therefore cannot fully simplify in the presence of metavariables (see Section 4.3). Furthermore, the technique supports function symbols with variable arity by interleaving an infinite number of searches when the function is applied, each assuming the function has a different arity.
The most common approach for ATP in proof assistants is called a hammer. In these systems, a goal is monomorphized into first-order logic [36, 14] which is then fed to a first-order theorem prover [31, 50, 46] or SMT solver [39, 5]. If the statement holds, a proof reconstruction tool builds a proof in the target language, often by re-proving the goal with a proof-producing first-order solver [26, 12]. This approach has been successful in languages that are built on higher-order logic [37, 28, 29] but has also been implemented in Rocq [16] and Metamath [11]. These tools are adept at solving goals in the first-order fragment of the target language.
@@ 4.1 Format
Note that with this definition, we have direct access to the head symbol of any term. The variables $\overline{x}$ and let declarations $\overline{y}$ are in a simultaneous declaration, such that they may depend on each other in any order, including circularly. The variable $f$ must be bound by a lambda binding or let definition in the term’s local context. A type in Canonical also has only one constructor, which is a term annotated with types for each of its bound variables:
@@
The types $\overline{X}$ and $\overline{Y}$ again may depend on any of the bound variables $\overline{x}$ and $\overline{y}$. For bound variable $f$ of type $\Pi\penalty 10000\ \overline{a}:\overline{Z}.\penalty 10000\ \texttt{let}\penalty 10000\ \overline{y^{\prime}}:\overline{Y^{\prime}}:=\overline{M^{\prime}}.\penalty 10000\ h\penalty 10000\ \overline{C}$, we say that the term in line 1 has the type in line 2 when:
@@ 4.2 Refinement
We associate with each metavariable a typing constraint. The starting metavariable’s type is constrained to be the goal type under definitional reduction. Consider the case of a metavariable ?M with the type in line 2 being refined by function symbol $f$ of type $\Pi\penalty 10000\ \overline{a}:\overline{Z}.\penalty 10000\ \texttt{let}\penalty 10000\ \overline{y^{\prime}}:\overline{Y^{\prime}}:=\overline{M^{\prime}}.\penalty 10000\ h\penalty 10000\ \overline{C}$. Let $\overline{\texttt{?A}}$ be the newly created metavariables to apply $f$. We generate the following equational constraint to match line 3:
@@
at which point the original typing constraint becomes inactive. If we find that one of our equational constraints is violated, we immediately backtrack the decision, reactivate the original typing constraint, and continue along another path. We do not need to backtrack the choices of metavariable, only the choices of head symbol, since a given term can be obtained by any ordering of metavariable refinement. The local context of a new metavariable $\texttt{?A}_{i}$ consists of the local context of ?M appended with the bound variables on the right hand side of the typing constraint in line 6.
@@ 4.3 Explicit Substitutions
cannot be reduced as it is not clear whether ?X contains $x$. It is essential that our algorithm be able to observe that this equation is violated since $f$ and $b$ are not equal, so that we may cut the branch and backtrack. Canonical depends on the ability to optimally determine whether an equational constraint has been violated, so that we only search over partial terms that are type correct.
@@
This operation can be performed in constant time with respect to the size of the terms and the number of arguments. Furthermore, $\beta$-reduction never creates new terms, only appending existing terms to existing explicit substitutions. If the first component of this term is a metavariable, $\beta$-reduction cannot continue and we say the reduction is stuck.
@@ 4.3.1 Equations
By default, Canonical considers such equations to be violated, but provides a program synthesis mode where they are considered stuck, allowing Canonical to synthesize functions that speculatively use recursors and projections even in situations where there are rigid equations. Additionally, Canonical is forbidden from using a constructor as the head symbol of the major argument of a recursor or projection, as the resulting term would not be in canonical form.
@@ 4.4 Search
We use iterative-deepening DFS so that refinement may be performed by directly mutating the assignment of a metavariable, rather than updating a persistent data structure. The cost of this is that we must retrace the steps of the prior levels each time we increase the entropy threshold. To mitigate this, we choose the threshold such that each level searches twice as many nodes as the previous. If this factor were smaller, there would be too much overhead from the redundant searching of prior levels. If it were larger, we might overshoot the number of steps necessary to find the proof term by a significant amount. There is a precise notion in which this factor of two is optimal.
@@ 4.4.1 Entropy
For the metavariables that have yet to be refined, we would like to estimate the number of refinements that it will take to completely solve them. Under the assumption that these metavariables are independent, the total refinement count to complete this branch will be the product of the individual refinement counts. To perform this estimate, we take advantage of the fact that our search has likely seen many similar metavariables in the past. While each metavariable may have a unique type and local context, we can bin metavariables into groups based on properties that have a small, finite number of combinations. Canonical bins metavariables according to whether they have a rigid equation, and their type ignoring the explicit substitution, representing the type in the original input program that this metavariable derives from. Each time a recursive DFS call returns, we accumulate search information about the metavariable being refined, such as the number of refinements performed on it or its children, whether the metavariable was successfully completed, or whether all branches resulted in an equation violation. After each iteration of the iterative-deepening search, we accumulate the statistics from all of the threads into the statistics used for entropy calculation.
@@ 4.4.2 Metavariable Ordering
Barring the rare case of an immediately terminating branch, the only way to stop the search is by increasing the entropy of the partial term as quickly as possible, to convince the DFS that the branch is unpromising. To do this, we prioritize metavariables with high entropy, as they likely are the most difficult to solve and harbor child metavariables with high entropy which accelerate the branch toward termination. If a metavariable has sufficiently high entropy, we prefer it over a metavariable that occurs later in the partial term but with less entropy.
@@ 5 Lean Tactic
The app constructor represents the application of a function expression to an argument expression. When translating this, we append the translation of the argument to args, and continue translation on the function. Once a constant symbol (const) or free variable (fvar) is reached, we output the final Term or Typ, with head being a string corresponding to that symbol.
@@
Canonical has no built-in object-theoretic notion of Sort. We add a let definition at the top level for “Sort” with no implementation and type “Sort.” This is similar to Lean’s judgment Sort u : Sort (u + 1), but removing the universe level annotation. In some circumstances this leads to unsoundness by Girard’s paradox, but in practice we have not seen Canonical solve a goal in this way. When we encounter a literal, we create a let declaration to represent it, with type “Nat” or “String.” Due to the size of literals, we generally do not include a definition for the let declaration, unless it is a natural number less than or equal to 5. In this case, we provide a definition in terms of the constructors of Nat.
@@ 5.2 Definitions
It is important to emphasize that the recursive translation of constant symbols and their definitions continues until primitive declarations are reached. This can violate abstraction barriers and lead to an explosion in the number of declarations if, for instance, real numbers are used. Deciding on which definitions to include is a question of relevance filtering and is beyond the scope of this work. We currently use a faithful recursive translation for completeness, but it is straightfoward to limit the definitions that Canonical receives.
@@ 6 Results
We have composed a Lean file with the 74 unique statements proven in the Natural Number Game. [^2] Some statements are duplicated to showcase the use of specialized tactics. The Natural Number Game uses a custom definition of natural numbers and the operations on them for pedagogical reasons. Our file instead uses the Lean standard library definition of Nat to allow fair comparison with tactics that have dedicated Nat logic, to more accurately represent real use cases, and to allow the use of definitional reduction. It is important to note that Canonical does not have specialization toward natural numbers or equality, and indeed (automatically) receives the inductive definition of these types separately with each problem.
@@
Figure 2: Natural Number Game problems solved by each tactic, separated by world.
Figure 2 shows the problems solved by each tactic, separated by world. Aesop solves 36% of the problems, Duper solves 61%, and Canonical solves 84%. This difference is primarily because Aesop and Duper are not intended to perform induction, and were not able to solve any problems requiring it. The first problem that Canonical fails on is in the Algorithm world:
@@ 8 Conclusion
Canonical is the first system capable of searching through terms in DTT, the foundations of Lean, Rocq, and Agda. The expressiveness of DTT and the generality of Canonical allows it to perform generalized induction on indexed inductive types, use higher-order axioms, and generate multiple constructive inhabitants of a type—all by exhaustive enumeration at the foundational level. We believe that search is a core technology that is useful for automated proof, program synthesis, counterexample generation, pedagogy, proof minimization, and data generation for AI. Beyond this, robust goal-directed reasoning is a necessary capability to reason at the human level.

### v2

@@ 1 Introduction
A powerful class of automation tools are those which translate a goal into first order logic (FOL) or higher order logic (HOL), invoke a solver like Vampire [32], and reconstruct the proof to close the goal. These are referred to as hammers [38, 16, 29, 30, 11]. By design, these systems can only use the built-in logical connectives, cannot directly reason with dependent types, and cannot synthesize functions or objects constructively. Hammers do not present an axiomatic proof term to the user, as such a term would be much too large to be read.
In this paper, we introduce Canonical, the first solver for type inhabitation in dependent type theory. Canonical searches exhaustively for terms of any type involving dependent products ($\Pi$-types), structure types, inductive types, and let definitions. The algorithm, discussed in Section 3, follows similarly to a theoretical account of sound and complete type inhabitation in DTT given by Dowek [19], proceeding via iterative refinement. In Section 4, we introduce a Lean tactic, canonical, which uses Canonical to find a proof term for the main goal. When it finds a proof term, it presents it to the user as a suggested replacement. While our implementation is not yet fully general, [^1] Specifically, we do not translate quotient types or universe levels; see Section 4. our goal is to approach soundness and completeness for the entirety of the Lean language, including definitional equality.
In Section 5, we demonstrate the performance of our tactic on the Natural Number Game for Lean 4 [10]. Canonical achieves an 84% on this benchmark, outperforming all existing Lean tactics. Canonical accomplishes this without any native support for natural numbers, propositional equality, logical connectives, or any other Lean definitions, and is instead provided the direct DTT representation at runtime. As such, these capabilities extend to user-defined datatypes and axiom schemas. We specifically highlight the higher-order reasoning and program synthesis capabilities of Canonical and show examples involving indexed inductive types and generalized induction.
@@ 2 Examples
Canonical is not limited to proving propositions. In this example we have an indexed inductive type Vec $\alpha$ n, representing vectors of length n with elements of type $\alpha$. Canonical is able to synthesize the append function using only the type signature, due to the enforced length constraints. The definition must be marked “noncomputable” in Lean because the code generator presently cannot compile recursors. If the return type Vec $\alpha$ (m + n) is changed to Vec $\alpha$ (n + m), Canonical will instead generate the function that appends the vectors in the reverse order.
@@
Since Canonical searches over inhabitants of a goal type, it can also be used to find multiple inhabitants. In this example, we use Canonical to find the two endomorphisms that exist over all groups – the identity map and the constant function which returns the identity element. Using the count flag, we obtain five inhabitants with two shown above. We use a simplified group definition, Group’, to avoid reasoning about auxiliary notions like exponents.
@@ 3.1 Format
Note that with this definition, we have direct access to the head symbol of any term. The variables $\overline{x}$ and let declarations $\overline{y}$ are in a simultaneous declaration, such that they may depend on each other in any order, including circularly. [^2] It is the responsibility of the encoding system to provide Canonical with a signature that is type-correct in the target language, such that this circularity is not inconsistent. The variable $f$ must be bound by a lambda binding or let definition in the term’s local context. A type in Canonical also has only one constructor, which is a term annotated with types for each of its bound variables:
@@
The types $\overline{X}$ and $\overline{Y}$ again may depend on any of the bound variables $\overline{x}$ and $\overline{y}$. This system is similar to the Logical Framework [24], although without universes. The let definitions are not essential for understanding the theory, but are used to encode constant symbols, reducible expressions, and let expressions from Lean. For bound variable $f$ of type $\Pi\penalty 10000\ \overline{a}:\overline{Z}.\penalty 10000\ \texttt{let}\penalty 10000\ \overline{y^{\prime}}:\overline{Y^{\prime}}:=\overline{M^{\prime}}.\penalty 10000\ h\penalty 10000\ \overline{C}$, we say that the term in line 1 has the type in line 2 when:
@@ 3.2 Refinement
We associate with each metavariable a typing constraint. The type of the starting metavariable is constrained to be definitionally equal to the goal type. Consider the case of a metavariable ?M with the type in line 2 being refined by function symbol $f$ of type $\Pi\penalty 10000\ \overline{a}:\overline{Z}.\penalty 10000\ \texttt{let}\penalty 10000\ \overline{y^{\prime}}:\overline{Y^{\prime}}:=\overline{M^{\prime}}.\penalty 10000\ h\penalty 10000\ \overline{C}$. Let $\overline{\texttt{?A}}$ be the newly created metavariables to apply $f$. We generate the following equational constraint to match line 3:
@@
at which point the original typing constraint becomes inactive. If we find that one of our equational constraints is violated, we immediately backtrack the refinement decision, reactivate the original typing constraint, and continue along another path. We do not need to backtrack the choices of metavariable, only the choices of head symbol, since a given term can be obtained by any ordering of metavariable refinement. The local context of a new metavariable $\texttt{?A}_{i}$ consists of the local context of ?M appended with the bound variables on the right hand side of the typing constraint in line 6.
@@ 3.3 Explicit Substitutions
cannot be reduced as it is not clear whether ?X contains $x$. It is essential that the algorithm be able to observe that this equation is violated since $f$ and $b$ are not equal, so that we may cut the branch and backtrack. Canonical depends on the ability to determine whether an equational constraint has been violated as early as is possible, so that we only search over partial terms that are type correct.
@@
This operation can be performed in constant time with respect to the size of the terms and the number of arguments. Furthermore, $\beta$-reduction never creates new terms, only appending existing terms to existing explicit substitutions. If the first component of this term is a metavariable, $\beta$-reduction cannot continue and we say the reduction is stuck. The let definitions from the input format encode nicely into this representation, simply extending the explicit substitution with each definition term.
@@ 3.3.1 Equations
By default, Canonical considers such equations to be violated, which is generally the desired behavior for type-directed problems like theorem proving. However, Canonical additionally provides a program synthesis mode where these equations are considered stuck, so it can synthesize functions that speculatively use recursors and projections even in situations where there are rigid equations. Additionally, Canonical is always forbidden from using a constructor as the head symbol of the major argument of a recursor or projection, as the resulting term would not be in canonical form.
@@ 3.4 Search
We use iterative-deepening DFS so that refinement may be performed by directly mutating the assignment of a metavariable, rather than updating a persistent data structure. The cost of this is that we must retrace the steps of the prior levels each time we increase the entropy threshold. To mitigate this, we choose the threshold such that each level searches twice as many nodes as the previous. If this factor were smaller, there would be too much overhead from the redundant searching of prior levels. If it were larger, we might overshoot the number of steps necessary to find the proof term by a significant amount. This factor of 2 minimizes the product of these effects.
@@ 3.4.1 Entropy
For the metavariables that have yet to be refined, we would like the entropy to be an estimate the number of refinements that it will take to completely solve them. Under the assumption that these metavariables are independent, the total refinement count to complete this branch will be the product of the individual refinement counts. To perform this estimate, we take advantage of the fact that our search has likely seen many similar metavariables in the past. While each metavariable may have a unique type and local context, we can bin metavariables into groups based on properties that have a small, finite number of combinations. Canonical bins metavariables according to whether they have a rigid equation, and their type ignoring the explicit substitution, representing the type in the original input program that this metavariable derives from. Each time a recursive DFS call returns, we accumulate search information about the metavariable being refined, such as the number of refinements performed on it or its children, whether the metavariable was successfully completed, or whether all branches resulted in an equation violation. After each iteration of the iterative-deepening search, we accumulate the statistics from all of the threads into the statistics used for entropy calculation.
@@ 3.4.2 Metavariable Ordering
Barring the rare case of an immediately terminating branch, the only way to stop the search is by increasing the entropy of the partial term as quickly as possible, to convince the DFS that the branch is unpromising. To do this, we prioritize metavariables with high entropy, as they likely are the most difficult to solve and harbor child metavariables with high entropy which accelerate the branch toward termination. If a metavariable has sufficiently high entropy, we prefer it over a metavariable that occurs later in the partial term but with less entropy. So, when selecting variables for refinement, we prefer selections that keep the entropy low, and when selecting the metavariable to refine, we prefer selections that increase the entropy as fast as possible.
@@ 4 Lean Tactic
The app constructor represents the application of a function expression to an argument expression. When translating (app $e_{1}\penalty 10000\ e_{2}$), we append the translation of the argument $e_{2}$ to args, and continue translation on the function expression $e_{1}$. Finally, once a constant symbol (const) or free variable (fvar) is reached, we set head to the name of the symbol, and return the complete Term.
@@
Canonical has no built-in object-theoretic notion of Sort. We add a let definition at the top level for Sort with no implementation and type Sort. This is similar to Lean’s judgment Sort u : Sort (u + 1), but removing the universe level annotation. In some circumstances this leads to unsoundness by Girard’s paradox, but in practice we have not seen Canonical solve a goal in this way. When we encounter a Nat or String literal, we create a let declaration to represent it, with the appropriate type. Due to the size of literals, we generally do not include a definition for the let declaration, unless it is a natural number less than or equal to 5. In this case, we provide a definition in terms of the constructors of Nat.
@@ 4.2 Definitions
It is important to emphasize that the recursive translation of constant symbols and their definitions continues until primitive declarations are reached. This can violate abstraction barriers and lead to an explosion in the number of declarations if, for instance, real numbers are used. Deciding on which definitions to include is a question of relevance filtering and is beyond the scope of this work. We currently use a faithful recursive translation for completeness, but it is straightforward to limit the definitions that Canonical receives.
@@ 5 Results
We have composed a Lean file with the 74 unique statements proven in the Natural Number Game. [^3] Some statements are duplicated to showcase the use of specialized tactics. The Natural Number Game uses a custom definition of natural numbers and the operations on them for pedagogical reasons. Our file instead uses the Lean standard library definition of Nat to allow fair comparison with tactics that have dedicated Nat logic, to more accurately represent real use cases, and to allow definitional reduction of some functions that the Natural Number Game defines axiomatically. It is important to note that Canonical does not have specialization toward natural numbers or equality, and indeed (automatically) receives the inductive definition of these types separately with each problem.
@@
Table 1: Natural Number Game problems solved by each tactic, separated by world.
@@
Table 1 shows the problems solved by each tactic, separated by world. Aesop solves 36% of the problems, Duper solves 61%, and Canonical solves 84%. This difference is primarily because Aesop and Duper are not intended to perform induction, and were not able to solve any problems requiring it. The first problem that Canonical fails on is in the Algorithm world:
@@
6 Related Work
There are a number of existing theorem proving systems that are based on type inhabitation. There are sauto [15] for Rocq, Agsy [35], Mimer [52], and Auto [31] for Agda, and TWELF [51] for the Logical Framework [24]. An essential subroutine for these systems is unification, in which two terms with existentially quantified variables are made equal to each other by a substitution on those variables. An algorithm for first-order unification was described by Robinson in 1965 [49]. In the higher-order setting, unification can have multiple (or infinitely many) solutions and is undecidable [25, 22]. Nonetheless, Huet provided a procedure to enumerate higher-order unifiers in 1975 [26]. To date, no theorem provers use this algorithm on account of its inefficiency, and instead elect for decidable and efficient pattern unification [39] (or extensions thereof, e.g. [2]). In particular, the restriction to pattern unification ensures that the solutions to unification problems are unique. While this handles many common cases, this means that these systems cannot perform general higher-order reasoning, such as the synthesis of induction predicates. Beyond this, these systems are not designed to handle the complexity of problems we are interested in solving with Canonical, such as those involving definitional equalities and requiring several nested recursors.
[…]
While these systems all have type theory support, they each sacrifice completeness or generality in some way. The only account of a complete search procedure for dependent type theory was given by Dowek [19]. Our algorithm is conceptually that of Dowek, although we use a custom explicit substitution representation for performance. To our knowledge, Dowek’s algorithm has not been implemented.
The most common approach for ATP in proof assistants is called a hammer. In these systems, a goal is monomorphized into first-order logic [37, 14] which is then fed to a first-order theorem prover [32, 54, 50] or SMT solver [41, 5]. If the statement holds, a proof reconstruction tool builds a proof in the target language, often by re-proving the goal with a proof-producing first-order solver [27, 12]. This approach has been successful in languages that are built on higher-order logic [38, 29, 30] but has also been implemented in Rocq [16] and Metamath [11]. These tools are adept at solving goals in the first-order fragment of the target language. There are also standalone tactics that invoke SMT solvers to solve first-order goals [40, 21].
@@ 8 Conclusion
Canonical is the first system capable of searching through terms in DTT, the foundations of Lean, Rocq, and Agda. The expressiveness of DTT and the generality of Canonical allows it to perform generalized induction on indexed inductive types, use higher-order axioms, and generate multiple constructive inhabitants of a type – all by exhaustive enumeration at the foundational level. We believe that search is a core technology that is useful for automated proof, program synthesis, witness/counterexample generation, pedagogy, proof minimization, and data generation for AI. Beyond this, robust goal-directed reasoning is a necessary capability to reason at the human level.